OpenStudy (anonymous):

i have this final and for the passed week we had a sub and i dont understand anything if someone is willing to help me even though its alot of questions i am willing to mail anyone five dollars for helping me i would offer more but thats all i have? my first question is 2x+3y+3z=5 4x+5y+3z=13 using matrices A.(5,13,13) B. (2,4,3) C. (1,2,-1) D. (1,3,-2) my second question is given the hyperbola (x2/4)-(y^2/9)=1,find asymptotes. A. x=(1/3)y,x=(-1/13)y B.y=(1/3)x,y=(-1/3)x C.y=(3/2)x,y+(-3/2)x D.x=(3/2)y,x=(-3/2)y

OpenStudy (anonymous):

my third what is the equation of the following graph A.x^2+y^2=1 B.x^2-y^2=1 C.x^2+y^2=2 D.x^2+y^2=-2 my fourth question solve the system by graphing y=-x+4(x-3)^2+y^2=25 A.(0,4) and (-3,7) B.(0,5) and (-3,7) C.(4,0) and (-3,7) D.(4,0) (7,-3) my fifth question use the factor theorem to determine (x-1) is a factor of x^3+x^2-16x-16? A. factor b. not factor my sixth solve using matrices 7x+5y=14 4x+3y=9 A.(7,-3) B.(-3,7) C. (3,-7) D.(7,3) My seventh question what is the 3rd term of (a-b)? A.21a^3b^4 B.21a^4b^3 C. 21a^3b^2 D.21a^6b My eighth question is what does the graph x^2-y^2=25 represent? A.circle B.ellipse C.hyperbola D. parabola my nineth question what is the system of equation describes the graph? my tenth question which of the following is the graph of the polynomial f(x)=(x-2)^3?  my eleventh question solve using matrices 7x+5y=14 4x+3y=9 A)(7,3) B)(-3,7) C)(3,-7) D)(7,3) My twelveth question is which of the following not a geometric sequence? A)1,2,4,8 B)3,1,-1,-3 C)1/2,1/4,1/8,1/16 D)216,72,24,8